Suppose that a liquid has an appreciable compresibility. Its density therefore varies with depth and pressure. The density at the surface is ρ0

Show that the density varies with pressure according to ρ=ρoekp   (Where p is gauge pressure at any depth, and k is the compressibillity.)

Find p as a function of depth, y.
Answer #1

According to Compressiblity definition, we know that:

k = \frac{1}{V}\frac{\Delta V}{\Delta P}\\ k \Delta P= \frac{\Delta V }{dV}

Integrating both side we get:

kP= \ln V\\ V=e^{kP}

Again we know that:

density=\frac{mass}{volume}\\ \rho =\frac{m}{V}

Put the value of V we get:

\rho =\frac{m}{e^{kP}}=\rho_0 e^{kP}

Where \rho_0 is constant.

From Pascal's Law, we know that:

\frac{dP}{dy} = \rho g\\ dP=\rho g dy

Integrating both side we get:

\int dP=\int \rho g dy\\ P=\rho g \int dy\\ P=\rho g y

So p is a function of depth ,y.

Proved.